# Integrability versus chaos in the steady state of many-body open quantum systems

### v1

@@ III.2 XX-chain with dephasing (model $B$)
Figure 3: Level statistics of $\mathcal{L}$ and $H_{ss}$ and NESS ETH in models $B$, $B1$, $B2$, and $B3$. (a) Complex spacing ratios of the Liouvillian $\mathcal{L}$ for the steady-state sector, with system size $N=9$ and dephasing strength $\gamma_{\mathrm{deph}}=1$. (b) and (c) Level spacing ratios of the steady-state Hamiltonian $H_{ss}$ for the zero magnetization sector, with system size $N=12$, for two values of dephasing strength, $\gamma^{\mathrm{deph}}=0.1\sim 1/N$ (b) and $\gamma^{\mathrm{deph}}=1\gg 1/N$ (c). (d)–(f) Standard deviation $\sigma$ of matrix elements of $S^{z}_{2}$ as a function of the system size, for $\gamma^{\mathrm{deph}}=0.1$ (d), $\gamma^{\mathrm{deph}}=1$ (e), and $\gamma^{\mathrm{deph}}=1/N$ (f). $\sigma$ is calculated in a window of width $0.5$ in the center of the spectrum.
@@
In model $B$, both the steady state and the Liouvillian are integrable. As expected, both objects have Poisson level statistics, see Fig. 3(a)–(b) for $\gamma^{\mathrm{deph}}=1$. Moreover, the steady state does not obey NESS ETH, see the blue curve in Fig. 3(d). We note that for small $\gamma^{\text{deph}}$ there is a finite-size effect to consider when comparing the signatures of the CSR with the expectation: The spectrum of the Liouvillian is then stretched along the real axis, as the relative strength of the dissipation gets smaller. For small system sizes, this leads to an increase in the probability of small angles $\theta$ between nearest- and next-to-nearest-neighbours. We find this represented in a deviation in the signature $\langle\cos\theta\rangle$ from the expectation for $\gamma^{\text{deph}}=0.1$ in model $B$ (see Table 2).
@@
We break integrability by considering next-to-nearest-neighbor and anisotropic interactions, resulting in models $B1$$B3$. We observe that in all three cases the Liouvillian follows GinUE statistics, reflecting their nonintegrability, see Fig. 3(a). Similarly, the steady state displays GUE level statistics, see Fig. 3(b). The variance of $S^{z}_{2}$ scales as $D^{-1/2}$ for large enough $N$, Fig. 3(d); the steady state thus obeys NESS ETH.
@@
We repeated the same calculations for another value of the dephasing strength, namely $\gamma^{\text{deph}}=1$, see Figs. 3(c) and (e). In this case, the steady state shows the characteristics associated with integrability: Poisson level statistics and the absence of NESS ETH. This goes against the intuition that models $B1$$B3$ are fully nonintegrable and consistently show the respective signatures for smaller $\gamma^{\text{deph}}$. However, note that in the limit of $\gamma^{\text{deph}}\gg\gamma_{l,r}^{\pm}$, dephasing dominates the dissipative dynamics and the steady state crosses over to a featureless infinite-temperature steady state. We conjecture that this crossover is perturbative and occurs on a scale $\gamma^{\text{deph}}\sim 1/N$, since the dephasing couples the system to the environment at every site of the chain, while the driving acts only on a finite number of sites (the boundaries). For the system sizes available, $\gamma^{\text{deph}}=1\gg 1/N$ and we are probing the trivial infinite-temperature state instead of the nontrivial boundary-driven NESS. A relation between the strength of dissipation and the support of the dissipator was observed in studies of the Liouvillian gap $g\sim\gamma N$ [98, 72]. A heuristic argument behind this scaling is as follows. If the dissipator is a sum of local superoperators acting only on one site $i$, $\mathcal{D}=\sum_{i}\mathcal{D}_{i}$, then $\mathcal{D}$ acts on an operator proportionally to both the dissipation strength and the support of the operator [70, 72]. Since in chaotic systems operators grow until they spread over the entire system, every operator will decay on a timescale of at most $\sim\gamma N$, which lower bounds the gap. Therefore, we conjecture that to reach a nontrivial steady state, we must set $\gamma^{\text{deph}}=\widetilde{\gamma}^{\text{deph}}/N$, with $\widetilde{\gamma}^{\text{deph}}$ constant. We show the results for this scaling of the dephasing strength in Fig. 3(f). For the system sizes available ($N\approx 10$), this results in $\gamma^{\text{deph}}\approx 1/10$ and we recover results similar to Fig. 3(d). Note that the bulk spectral statistics of the Liouvillian are unaffected by large $\gamma^{\mathrm{deph}}$ (while $\rho_{ss}$ becomes trivial if $N\gamma^{\mathrm{deph}}\gg\gamma^{\pm}_{l,r}$, $\mathcal{L}$ does not). This can be seen in Fig. 3(a), where we show the complex level spacings for $\gamma^{\mathrm{deph}}=1\gg 1/N$, which perfectly conform to non-Hermitian RMT.
IV Structure of the steady-state Hamiltonian $H_{ss}$
@@
Having determined the (non-)integrability of the steady states of the different models in the previous section through the use of spectral statistics, we now turn to the investigation of the structure of $H_{ss}$ in terms of the operators contained in it. A natural assumption would be that integrable steady-state Hamiltonians are “simple” in some sense, e.g., they are linear combinations of local few-body operators, perhaps closely related to known integrable Hamiltonian models. In this section, we will investigate the operator expansion of $H_{ss}$ and show that this is not the case: even integrable $H_{ss}$ have contributions from all possible operators (including many-body and nonlocal ones). Steady-state integrability is therefore conceptually distinct from Hamiltonian and Liouvillian integrability. Nevertheless, we shall show that the weight of different contributions to the operator expansion of $H_{ss}$ gives a sharp discriminator between chaotic and integrable steady states.
@@ IV.1 Operator size and size distribution
We expand $H_{ss}$ in the basis of Pauli strings to motivate a measure for its size and in particular the contributions from few-body operators [67, 71, 70, 68, 72, 69]. A Pauli string $\hat{R}$ is any product $\prod_{j=1}^{N}\sigma_{j}^{\alpha}$ where $\alpha=0,x,y,z$, $\sigma^{0}=\mathds{1}$ and $\sigma^{x,y,z}$ are the standard $2\times 2$ Pauli matrices. There are $4^{N}$ distinct Pauli strings, which form an orthonormal basis of the space of operators acting on $\mathcal{H}$. An operator is expanded in terms of Pauli strings as
@@
Figure 4: Size distribution $P(S)$ of $H_{ss}$, given by Eq. (21), for different system sizes $N$ and models of type $A$ (top) and $B$ (bottom). Top: Blue curves correspond to model $A$ and red curves to model $A^{\prime}$. Bottom: Green curves correspond to a dephasing strength $\gamma^{\text{deph}}=1/N$ and orange curves to $\gamma^{\text{deph}}=1$. The size distributions for $B3$ are not shown, as they are very similar to those of $B2$.
@@ IV.2 Coefficients for short Pauli strings
Figure 5: Coefficients of length-one Pauli strings $\sigma_{i}^{z}$, depending on their position $i$ in the chain, in $H_{ss}$ of models of type $A$ and $B$. Top (type $A$): Red curves correspond to $H_{ss}$ of model $A$ and blue curves to $H_{ss}$ of model $A^{\prime}$. Bottom (type $B$): Green curves correspond to a dephasing strength $\gamma^{\text{deph}}=1/N$ and orange curves to $\gamma^{\text{deph}}=1$. The coefficients for $B3$ are not shown as they are very similar to those of $B2$.
Having determined that strings of all lengths contribute to $H_{ss}$ but that shorter strings have larger contributions, in this section, we investigate the structure of the coefficients $c_{\hat{R}}$ for the Pauli strings $\hat{R}$ of lengths one and two.
As $H_{ss}$ commutes with $S^{z}$, we can immediately write down the following restrictions on the coefficients:
@@ V Summary and Discussion
We verified that the integrability of the steady state is signaled by Poissonian level statistics for the steady-state spectrum and a violation of NESS ETH, irrespective of whether or not the full Liouvillian is integrable. (We used the spectrum of the “steady-state Hamiltonian” $H_{ss}=-\ln\rho_{ss}$ rather than that of $\rho_{ss}$ itself, and have checked that this does not make any difference.)
Integrability of the steady state raises the question of whether, for known examples of integrable steady states, the object $H_{ss}$ has especially simple properties. Do we obtain known integrable Hamiltonians? Do we at least obtain Hamiltonians with few-body local operators? For the known cases, we have addressed this question (Sec. IV) and the answer is definitely negative. As Fig. 4 demonstrates, operators of all lengths contribute for integrable models and, unsurprisingly, also for nonintegrable models. Figure 6 shows that even the two-body operators are not particularly well-behaved — they are generally long-range. However, we also found that there are noticeable differences with the nonintegrable cases: in general, the operator size distribution for integrable models is biased more strongly toward shorter-length operators. In fact, in Fig. 4 the integrable cases can be visually distinguished.
@@
In general, the results of this work described in Sec. III confirm what one might have expected from the two types of integrability. The only subtlety was the appropriate strength of the dephasing coupling. The results of Sec. IV were less predictable, and nontrivial open questions remain. It seems likely that integrability of the NESS should render $H_{ss}$ to be ‘simple’ in some manner, however, we have not found an obvious characterization of simplicity for the cases examined. A general understanding of what differences to expect for $H_{ss}$, between integrable and nonintegrable steady states, appears to be currently lacking and deserves further investigation.
@@ Acknowledgements.
LS was supported by a Research Fellowship from the Royal Commission for the Exhibition of 1851. MH acknowledges financial support from the Deutsche Forschungsgemeinschaft under grant SFB 1143 (project-id 247310070).

### v2

@@ III.2 XX chain with dephasing (model $B$)
Figure 3: Level statistics of $\mathcal{L}$ and $H_{\mathrm{ss}}$ and NESS ETH in models $B$, $B1$, $B2$, and $B3$. (a) Complex spacing ratios of the Liouvillian $\mathcal{L}$ for the steady-state sector, with system size $N=9$ and dephasing strength $\gamma_{\mathrm{deph}}=1$. (b) and (c) Level spacing ratios of the steady-state Hamiltonian $H_{\mathrm{ss}}$ for the zero magnetization sector, with system size $N=12$, for two values of dephasing strength, $\gamma^{\mathrm{deph}}=0.1\sim 1/N$ (b) and $\gamma^{\mathrm{deph}}=1\gg 1/N$ (c). (d)–(f) Standard deviation $\sigma$ of matrix elements of $S^{z}_{2}$ as a function of the system size, for $\gamma^{\mathrm{deph}}=0.1$ (d), $\gamma^{\mathrm{deph}}=1$ (e), and $\gamma^{\mathrm{deph}}=1/N$ (f). $\sigma$ is calculated in a window of width $0.5$ in the center of the spectrum.
@@
In model $B$, both the steady state and the Liouvillian are integrable. As expected, both objects have Poisson level statistics; see Fig. 3(a)–(b) for $\gamma^{\mathrm{deph}}=1$. Moreover, the steady state does not obey NESS ETH; see the blue curve in Fig. 3(d). We note that for small $\gamma^{\text{deph}}$ there is a finite-size effect to consider when comparing the signatures of the CSR with the expectation: The spectrum of the Liouvillian is then stretched along the real axis, as the relative strength of the dissipation gets smaller. For small system sizes, this leads to an increase in the probability of small angles $\theta$ between nearest- and next-to-nearest-neighbors. We find this represented in a deviation in the signature $\langle\cos\theta\rangle$ from the expectation for $\gamma^{\text{deph}}=0.1$ in model $B$ (see Table 2).
@@
We break integrability by considering next-to-nearest-neighbor and anisotropic interactions, resulting in models $B1$$B3$. We observe that in all three cases the Liouvillian follows GinUE statistics, reflecting their nonintegrability, see Fig. 3(a). Similarly, the steady state displays GUE level statistics; see Fig. 3(b). The variance of $S^{z}_{2}$ scales as $D^{-1/2}$ for large enough $N$, Fig. 3(d); the steady state thus obeys NESS ETH.
@@
We repeated the same calculations for another value of the dephasing strength, namely $\gamma^{\text{deph}}=1$; see Figs. 3(c) and (e). In this case, the steady state shows the characteristics associated with integrability: Poisson level statistics and the absence of NESS ETH. This goes against the intuition that models $B1$$B3$ are fully nonintegrable and consistently show the respective signatures for smaller $\gamma^{\text{deph}}$. However, note that in the limit of $\gamma^{\text{deph}}\gg\gamma_{l,r}^{\pm}$, dephasing dominates the dissipative dynamics and the steady state crosses over to a featureless infinite-temperature steady state. We conjecture that this crossover is perturbative and occurs on a scale $\gamma^{\text{deph}}\sim 1/N$, since the dephasing couples the system to the environment at every site of the chain, while the driving acts only on a finite number of sites (the boundaries). For the system sizes available, $\gamma^{\text{deph}}=1\gg 1/N$ and we are probing the trivial infinite-temperature state instead of the nontrivial boundary-driven NESS. A relation between the strength of dissipation and the support of the dissipator was observed in studies of the Liouvillian gap $g\sim\gamma N$ [103, 104, 76]. A heuristic argument behind this scaling is as follows. If the dissipator is a sum of local superoperators acting only on one site $i$, $\mathcal{D}=\sum_{i}\mathcal{D}_{i}$, then $\mathcal{D}$ acts on an operator proportionally to both the dissipation strength and the support of the operator [73, 76]. Since in chaotic systems operators grow until they spread over the entire system, every operator will decay on a timescale of at most $\sim\gamma N$, which lower bounds the gap. Therefore, we conjecture that to reach a nontrivial steady state, we must set $\gamma^{\text{deph}}=\widetilde{\gamma}^{\text{deph}}/N$, with $\widetilde{\gamma}^{\text{deph}}$ constant. We show the results for this scaling of the dephasing strength in Fig. 3(f). For the system sizes available ($N\approx 10$), this results in $\gamma^{\text{deph}}\approx 1/10$ and we recover results similar to Fig. 3(d). Note that the bulk spectral statistics of the Liouvillian are unaffected by large $\gamma^{\mathrm{deph}}$ (while $\rho_{\mathrm{ss}}$ becomes trivial if $N\gamma^{\mathrm{deph}}\gg\gamma^{\pm}_{l,r}$, $\mathcal{L}$ does not). This can be seen in Fig. 3(a), where we show the complex spacing ratios for $\gamma^{\mathrm{deph}}=1\gg 1/N$, which perfectly conform to non-Hermitian RMT.
IV Structure of the steady-state Hamiltonian $H_{\mathrm{ss}}$
@@
Having determined the (non)integrability of the steady states of the different models in the previous section through the use of spectral statistics, we now turn to the investigation of the structure of $H_{\mathrm{ss}}$ in terms of the operators contained in it. A natural assumption would be that integrable steady-state Hamiltonians are “simple” in some sense, e.g., they are linear combinations of local few-body operators, perhaps closely related to known integrable Hamiltonian models. In this section, we will investigate the operator expansion of $H_{\mathrm{ss}}$ and show that this is not the case: even integrable $H_{\mathrm{ss}}$ have contributions from all possible operators (including many-body and nonlocal ones). Steady-state integrability is therefore conceptually distinct from Hamiltonian and Liouvillian integrability. Nevertheless, we shall show that the weight of different contributions to the operator expansion of $H_{\mathrm{ss}}$ gives a sharp discriminator between chaotic and integrable steady states.
@@ IV.1 Operator size and size distribution
We expand $H_{\mathrm{ss}}$ in the basis of Pauli strings to motivate a measure for its size and in particular the contributions from few-body operators [70, 74, 73, 71, 75, 76, 72]. A Pauli string $\hat{R}$ is any product $\prod_{j=1}^{N}\sigma_{j}^{\alpha}$ where $\alpha=0,x,y,z$, $\sigma^{0}=\mathds{1}$ and $\sigma^{x,y,z}$ are the standard $2\times 2$ Pauli matrices. There are $4^{N}$ distinct Pauli strings, which form an orthonormal basis of the space of operators acting on $\mathcal{H}$. An operator is expanded in terms of Pauli strings as
@@
Figure 4: Size distribution $P(S)$ of $H_{\mathrm{ss}}$, given by Eq. (21), for different system sizes $N$ and models of type $A$ (top) and $B$ (bottom). Top: Blue curves correspond to model $A$ and red curves to model $A^{\prime}$. Bottom: Green curves correspond to a dephasing strength $\gamma^{\text{deph}}=1/N$ and orange curves to $\gamma^{\text{deph}}=1$. The size distributions for $B3$ are not shown, as they are very similar to those of $B2$.
@@ IV.2 Coefficients for short Pauli strings
Figure 5: Coefficients of length-one Pauli strings $\sigma_{i}^{z}$, depending on their position $i$ in the chain, in $H_{\mathrm{ss}}$ of models of type $A$ and $B$. Top (type $A$): Red curves correspond to $H_{\mathrm{ss}}$ of model $A$ and blue curves to $H_{\mathrm{ss}}$ of model $A^{\prime}$. Bottom (type $B$): Green curves correspond to a dephasing strength $\gamma^{\text{deph}}=1/N$ and orange curves to $\gamma^{\text{deph}}=1$. The coefficients for $B3$ are not shown as they are very similar to those of $B2$.
Having determined that strings of all lengths contribute to $H_{\mathrm{ss}}$ but that shorter strings have larger contributions, in this section, we investigate the structure of the coefficients $c_{\hat{R}}$ for the Pauli strings $\hat{R}$ of lengths one and two.
As $H_{\mathrm{ss}}$ commutes with $S^{z}$, we can immediately write down the following restrictions on the coefficients:
@@ V Summary and Discussion
We verified that the integrability of the steady state is signaled by Poissonian level statistics for the steady-state spectrum and a violation of NESS ETH, irrespective of whether or not the full Liouvillian is integrable. (We used the spectrum of the “steady-state Hamiltonian” $H_{\mathrm{ss}}=-\ln\rho_{\mathrm{ss}}$ rather than that of $\rho_{\mathrm{ss}}$ itself, and have checked that this does not make any difference.)
Integrability of the steady state raises the question of whether, for known examples of integrable steady states, the object $H_{\mathrm{ss}}$ has especially simple properties. Do we obtain known integrable Hamiltonians? Do we at least obtain Hamiltonians with few-body local operators? For the known cases, we have addressed this question (Sec. IV) and the answer is definitely negative. As Fig. 4 demonstrates, operators of all lengths contribute for integrable models and, unsurprisingly, also for nonintegrable models. Figure 6 shows that even the two-body operators are not particularly well-behaved — they are generally long-range. However, we also found that there are noticeable differences with the nonintegrable cases: in general, the operator size distribution for integrable models is biased more strongly toward shorter-length operators. In fact, in Fig. 4 the integrable cases can be visually distinguished.
@@
In general, the results of this work described in Sec. III confirm what one might have expected from the two types of integrability. The only subtlety was the appropriate strength of the dephasing coupling. The results of Sec. IV were less predictable, and nontrivial open questions remain. It seems likely that integrability of the NESS should render $H_{\mathrm{ss}}$ to be ‘simple’ in some manner, however, we have not found an obvious characterization of simplicity for the cases examined. A general understanding of what differences to expect for $H_{\mathrm{ss}}$, between integrable and nonintegrable steady states, appears to be currently lacking and deserves further investigation.
@@ Acknowledgements.
We thank M. Žnidarič for useful comments. LS was supported by a Research Fellowship from the Royal Commission for the Exhibition of 1851. MH acknowledges support from the Deutsche Forschungsgemeinschaft under grant SFB 1143 (project-id 247310070).